OpenStudy (anonymous):

evaluate the limit. estimate the value of lim x is approaching 0 (x)/√(1+3x)-1. b. make a table of values of f(x) for x close to 0 and guess the value of the limit.

OpenStudy (anonymous):

please help me!!!!

OpenStudy (anonymous):

is this \[\frac{x}{\sqrt{1+3x}-1}\]?

OpenStudy (anonymous):

limit is \[\frac{2}{3}\] but you are supposed to guess from the evidence, so get a calculator (or better yet a spread sheet) and start plugging in smaller and smaller numbers, like say .1, .01, .001 and see what you get

OpenStudy (anonymous):

yes that's it! how do you figure out the limit because i can't find my graphing calculator. so how did you find 2/3

OpenStudy (anonymous):

i know some easy tricks, like l'hopital's rule, but you don't know them yet. soon though...